OpenStudy (study_buddy99):

A youth group and their leaders visited mammoth cav. 2 adults and 5 students in one van paid $77 for a tour of the cave. Two adults and 7 students in a second van paid $95 for the same tour. Find the adult price and the student price of the ticket

OpenStudy (anonymous):

two more students cost how much more money?

OpenStudy (anonymous):

or in simple english, what is \(95-77\)?

OpenStudy (study_buddy99):

but that still makes no sense regarding what the question is asking

OpenStudy (anonymous):

why not?

OpenStudy (anonymous):

two students cost an extra \(95-77=18\) right?

OpenStudy (anonymous):

Okay lets think of our unknowns as variables. Lets just easily calls those variables x & y. What we don't know is how much adults paid and how much a student paid. What we do know is their cost. I can tell you our goal here is to solve a two linear equations. I'll help you find the first equation from the first part of the question, and i bet you'll be able to get the second one. \[2x+5y=77\]

OpenStudy (anonymous):

if two students cost an extra $18, the price per student is half of that

OpenStudy (study_buddy99):

but it's still asking the price for the student and adult tickets

OpenStudy (anonymous):

right two students cost $18 what is the price of one student?

OpenStudy (study_buddy99):

one student would be $9

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and since one student cost $9, then we can figure out what the adult cost since \[2\times \text{ adult price}+5\times \$9=\$77\]

OpenStudy (anonymous):

of if you prefer \[2\times \text{ adult price}+7\times \$9=\$95\] either way

OpenStudy (anonymous):

you know how to solve those?

OpenStudy (study_buddy99):

I see, thank you

OpenStudy (anonymous):

yw